Let f be the function given by f (x) = |x|. Which of the following statements about f is true?


I. f is continuous at x = 0

II. f is differentiable at x = 0

III. f has an absolute maximum at x = 0.


1
Expert's answer
2021-02-24T07:48:06-0500

Ans:-

f(x) "\Rightarrow" "\begin{cases}
x&x\ge0\\-x&x<0\end{cases}"

At "x<0" , f(x)=-x is continuous every where

At "x>0" , f(x)=x is continuous every where

at x=0

L.H.L.

"lim_{x \to o^-} f(x)=" "lim_{x \to o} -x=" 0

R.H.L.

"lim_{x \to o^+} f(x)=" "lim_{x \to 0} x="0

f(0)=0

L.H.L=R.H.L=f(0)

f(x) is continuous at x=0


Now,

f'(x)"\Rightarrow" "\begin{cases}
-1&x<0\\1&x\ge0\end{cases}"

At x=0, f'(x) is not differentiable


f''(x)=0 then

So f(x) is neither maximum nor minimum at x=0


Hence Given statement is false and first option is true which is f(x) is continuous at x=0